An analyst is using exponential smoothing to forecast the daily demand for a key product. The analyst starts with a naive forecast for time period 2, then begins using exponential smoothing with a smoothing constant of 0.15. The table below shows some of the calculations. period actual forecast 1 125 2 136 125 3 144 126.65 4 157 129.25 5 181 ? What is the predicted demand for time period 5? Round your answer to two decimal places.

Period Actual Forecast
1 125
2 136 125
3 144 126.65
4 157 129.25
5 181 ?
Answer #1

The predicted demand for the time

actual is 181 for 5th period and smoothing constant is 0.15

forecast demand FT = FT-1 +$(AT-1 - FT-1 )

WHERE FT = NEW FORECAST DEMAND

  FT-1 = PREVIOUS PERIOD FORECAST

$ = SMOOTHING CONSTANT

AT-1 = PREVIOUS PERIOD ACTUAL DEMAND

      FT-1 = PREVIOUS PERIOD FORECAST

based in these = 129.25+0.5(157-129.25)

= 129.75 + 27.75

= 157.50

SO THE PREDICTED DEMAND FOR PERIOD 5 IS 157.50